OpenStudy (princess_dymie):

how many hours are in 90,000 seconds? A)25h B)750h C)10,800,000h D)324,000,000h

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Recall there are 60 seconds in a minute and 60 minutes in an hour.

OpenStudy (skullpatrol):

Step 1 change 90, 000 seconds into minutes.

OpenStudy (avirllavinge2017):

First before I explain , what do you think? I already have answer determined in my head , what are your thoughts?

OpenStudy (skullpatrol):

The fact that there are 60 seconds in 1 minute can be stated as an equation as 60 sec = 1 min. This equation can be used as a conversion factor as $$\dfrac{60sec}{1min}$$ or $$\dfrac{1min}{60sec}$$ both of which equal 1. $$\huge 90,000sec \times \dfrac{1min}{60sec}=?$$